# ZSDEVC: Zero-Shot Diffusion-based Emotional Voice Conversion with Disentangled Mechanism

### v4

@@ 1 Introduction
Human speech is more than just a medium for words; it conveys an audible declaration of one’s identity and emotion [1]. Modifying emotional expressions while preserving linguistic content and speaker identity is the focus of emotional voice conversion (EVC), a technology that enhances user experiences in applications of human-machine interactions, virtual assistants, and entertainment industries [2, 3]. Advances in deep learning are the driving forces in the progression of EVC technology, which has been investigated mainly in the context of speaker-dependent emotion conversion, demonstrating cases of realistically expressive and natural-sounding synthesized speech [4, 5, 6, 7].
Many deep learning methods have been investigated to achieve emotional voice conversion. These methods fall mainly into two categories: adversarial generative networks (GAN) [4, 5] and autoencoders [6, 7]. GAN-based approaches use adversarial mechanisms to learn direct mappings between data distributions of different emotional states, allowing direct emotion conversion. In contrast, autoencoder-based methods decompose speech into distinct representation units, such as linguistic content, speaker identity, and emotional information, providing better control over emotion conversion. However, despite their advances, these techniques remain suboptimal in accurately converting emotional states and can introduce distortions in the converted voice, compromising its naturalness and overall quality. The application is further limited due to the data requirements of having well-crafted target speakers’ emotional speech samples.
@@ 2.1 Proposed Method
This work aims to solve the problem of zero-shot emotional voice conversions. Given a pair of $X_{src}:=g(c_{src},spk_{src},emo_{src})$ and reference $X_{ref}:=g(c_{ref},spk_{ref},emo_{ref})$ speech utterances, where each utterance is composed of linguistic content $c$, speaker identity $spk$, and emotion information $emo$ and $g(.)$ is a generative process, our proposed method $G$ aims to perform the conversion process $\hat{X}=G(c_{src},spk_{src},emo_{ref})$ that preserves both content and speaker identity while transforming emotion from $emo_{src}$ to $emo_{ref}$. We focus on a zero-shot scenario in which both the source and reference speech, as well as the speaker’s identity, remain entirely unseen during training.
@@ 2.1.1 Encoders
Three pre-trained encoders are used to capture linguistic content representations $c$, speaker identity $spk$, and emotional expression $emo$.
Phoneme Encoding: To encode linguistic content $\bar{X}$, we adapt a pre-trained transformer-based encoder from [16] to convert input mel-spectrograms $X_{0}$ into speaker and emotion independent “average-voice” mel features that replace each phoneme-level mel feature with the corresponding average phoneme-level mel features.
@@ 2.1.2 Diffusion Decoder
We employ the diffusion framework based on stochastic differential equations (SDE) described in [11], conditioned on given representations $\bar{X},Z_{s},Z_{e}$ to generate high-quality speech. The diffusion process gradually transforms the real sample ${X_{0}}$ into ${X_{t}}$ with time-step $t\in[0,1]$ that terminates at average-voice mel-spectrogram $\bar{X}$ when $t=1$ by adding Gaussian noise in a forward process; and generates ${X_{0}}$ from $\bar{X}$ by removing the corresponding score estimation $s_{\theta}(X_{t},t,\bar{X},\hat{Z_{s}},\hat{Z_{e}})$ in a reverse process. The $s_{\theta}$ with parameter $\theta$ is trained by minimizing mean square error loss $\mathcal{L}_{diff}$ between added noise and $s_{\theta}$.
@@ 2.1.4 Disentangled Loss
In addition to $L_{diff}$ for training diffusion-based decoder, we follow [13] to use a mel-spectrogram recontruction loss $\mathcal{L}_{rec}$ that measures the $\mathcal{L}_{1-norm}$ between $X_{0}$ and $\hat{X}_{0}$, where $\hat{X}_{0}$ is the single-step approximation relying on $X_{t},\bar{X},s_{\theta}$ using Tweedie’s formula [23]. We use $\lambda_{rec}=(1-t^{2})$ adapted from [13] to reduce the importance of the loss as $X_{t}$ becomes increasingly noisy due to added Gaussian noise at larger values of t.
@@
Target 3.606 0.816 3.429 3.155 1.000 4.197$\pm$0.178 4.402$\pm$0.150 0.889
StarGAN-EVC [5] Seen Speaker 3.128 0.884 3.461 3.190 0.222 4.000$\pm$0.188 3.863$\pm$0.228 0.299
Seq2Seq-EVC [6] Seen Speaker 1.903 0.663 3.301 2.957 0.444 1.872$\pm$0.177 2.872$\pm$0.236 0.120
Emovox [7] Seen Speaker 2.381 0.698 3.234 2.930 0.333 2.197$\pm$0.188 2.974$\pm$0.211 0.333
Prosody2Vec [9] Seen Speaker 2.482 0.730 3.071 2.717 0.889 2.803$\pm$0.221 3.308$\pm$0.237 0.769
EMOCONV-DIFF [13] Unseen Speaker 3.973 0.834 3.611 3.347 0.667 4.709$\pm$0.098 4.291$\pm$0.179 0.256
ZSDEVC (Proposed) Unseen Speaker 3.583 0.768 3.589 3.336 0.889 4.342$\pm$0.156 3.752$\pm$0.201 0.530
@@ 3.1.2 Evaluation Setup
To assess the effectiveness of our method, we then evaluated our methods in zero-shot scenarios on both in-the-wild datasets, MSP-Podcast, with real-world scenarios, and the act-out dataset, ESD, with high-quality recordings. We randomly sample 300 utterances of each emotion category with unseen speakers from both datasets to conduct the following experiments as source speech. We then perform zero-shot emotional voice conversions that include all the transformations between angry, happy, sad, and neutral, except transforming from emotional speech to neutral. We compared the methods under different training schemes and structures, i.e., using only $\mathcal{L}_{diff}+\mathcal{\lambda}_{rec}\mathcal{L}_{rec}$, which is the reimplementation of EMOCONV-DIFF, and using $\mathcal{L}_{Total}$ in equation 2 with additional layers for disentanglement. We then apply the proposed expressive guidance method on the model trained with $\mathcal{L}_{Total}$. We compared $EG_{spk}$, $EG_{emo}$, and $EG_{spk,emo}$ with different settings of negative condition that replace the representation of positive condition corresponding to speaker identity $spk$, emotion information $emo$, or both, respectively. We also evaluate the Source (VO), which represents the source speech reconstructed by the vocoder.

### v5

@@ 1 Introduction
Human speech is more than just a medium for words; it conveys an audible declaration of one's identity and emotion [1]. Modifying emotional expressions while preserving linguistic content and speaker identity is the focus of emotional voice conversion (EVC), a technology that enhances user experiences in applications of human-machine interactions, virtual assistants, and entertainment industries [2, 3]. Advances in deep learning are the driving forces in the progression of EVC technology, which has been investigated mainly in the context of speaker-dependent emotion conversion, demonstrating cases of realistically expressive and natural-sounding synthesized speech [4, 5, 6, 7].
Many deep learning methods have been investigated to achieve emotional voice conversion. These methods fall mainly into two categories: adversarial generative networks (GAN) [4, 5] and autoencoders [6, 7]. GAN-based approaches use adversarial mechanisms to learn direct mappings between data distributions of different emotional states, allowing direct emotion conversion. In contrast, autoencoder-based methods decompose speech into distinct representation units, such as linguistic content, speaker identity, and emotional information, providing better control over emotion conversion. However, despite their advances, these techniques remain suboptimal in accurately converting emotional states and can introduce distortions in the converted voice, compromising its naturalness and overall quality. The application is further limited due to the data requirements of having well-crafted target speakers' emotional speech samples.
@@ 2.1 Proposed Method
This work aims to solve the problem of zero-shot emotional voice conversions. Given a pair of $X_{src}:=g(c_{src},spk_{src},emo_{src})$ and reference $X_{ref}:=g(c_{ref},spk_{ref},emo_{ref})$ speech utterances, where each utterance is composed of linguistic content $c$, speaker identity spkspk, and emotion information emoemo and $g(.)$ is a generative process, our proposed method $G$ aims to perform the conversion process $\hat{X}=G(c_{src},spk_{src},emo_{ref})$ that preserves both content and speaker identity while transforming emotion from $emo_{src}$ to $emo_{ref}$. We focus on a zero-shot scenario in which both the source and reference speech, as well as the speaker's identity, remain entirely unseen during training.
@@ 2.1.1 Encoders
Three pre-trained encoders are used to capture linguistic content representations $c$, speaker identity spkspk, and emotional expression emoemo.
Phoneme Encoding: To encode linguistic content $\bar{X}$, we adapt a pre-trained transformer-based encoder from [16] to convert input mel-spectrograms $X_{0}$ into speaker and emotion independent ``average-voice" mel features that replace each phoneme-level mel feature with the corresponding average phoneme-level mel features.
@@ 2.1.2 Diffusion Decoder
[…]
@@ 2.1.4 Disentangled Loss
In addition to $L_{diff}$ for training diffusion-based decoder, we follow [13] to use a mel-spectrogram recontruction loss $\mathcal{L}_{rec}$ that measures the $\mathcal{L}_{1-norm}$ between $X_{0}$ and $\hat{X}_{0}$, where $\hat{X}_{0}$ is the single-step approximation relying on $X_{t},\bar{X},s_{\theta}$ using Tweedie's formula [23]. We use $\lambda_{rec}=(1-t^{2})$ adapted from [13] to reduce the importance of the loss as $X_{t}$ becomes increasingly noisy due to added Gaussian noise at larger values of t.
@@
[…]
@@ 3.1.2 Evaluation Setup
To assess the effectiveness of our method, we then evaluated our methods in zero-shot scenarios on both in-the-wild datasets, MSP-Podcast, with real-world scenarios, and the act-out dataset, ESD, with high-quality recordings. We randomly sample 300 utterances of each emotion category with unseen speakers from both datasets to conduct the following experiments as source speech. We then perform zero-shot emotional voice conversions that include all the transformations between angry, happy, sad, and neutral, except transforming from emotional speech to neutral. We compared the methods under different training schemes and structures, i.e., using only $\mathcal{L}_{diff}+\mathcal{\lambda}_{rec}\mathcal{L}_{rec}$, which is the reimplementation of EMOCONV-DIFF, and using $\mathcal{L}_{Total}$ in equation 2 with additional layers for disentanglement. We then apply the proposed expressive guidance method on the model trained with $\mathcal{L}_{Total}$. We compared $EG_{spk}$, $EG_{emo}$, and $EG_{spk,emo}$ with different settings of negative condition that replace the representation of positive condition corresponding to speaker identity spkspk, emotion information emoemo, or both, respectively. We also evaluate the Source (VO), which represents the source speech reconstructed by the vocoder.